OpenStudy (fisherdevo):

Apply the distributive property to factor out the greatest common factor. 44 + 48 = will fan and medal

OpenStudy (thatonegirl_):

I was trying to help, but I guess you don't want my help. Good luck.

OpenStudy (fisherdevo):

thanks

OpenStudy (abdullahm):

We need to find the greatest common factor. Do you know how to do that?

OpenStudy (abdullahm):

Here's how to find the greatest common factor. I'll do an example: 22, list the factors: 1, 2, 11 ,22 24, list the factors: 1, 2, 3, 4, 6, 8, 12, 24 the greatest common factor is 2 that is what we will factor out of 22 and 24 22 + 24 = (2 * 11) + (2 * 12) = 2(11 + 12)